# Random Scaling of Emergent Capabilities

### v4

@@ Appendix B Breakthroughness and Linearity
Srivastava et al. [2023] introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})},$
$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$
@@
In Figure 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and addition (below).
@@
Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 10(a) count and 10(b) addition.
Appendix C Additional Exact Match Accuracy Histogram Plots
@@
Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling 11(a) width and 11(b) depth.
@@
Our plots in the main paper fixed test length 40 for addition (Figure 2). In Figure 11 we show analogous histograms for a fixed test length of 60 on the count task. We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
In Section 2.5, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task. For example, Figure 12 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
@@ Appendix D Additional Continuous Metric Plots
In Section 2.4 we have seen seen that the addition histograms are bimodal in exact match accuracy (Figure 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Figure 4). We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
@@
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Figures 14 and 15 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
In Figure 16, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 11 and 15 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
Finally, Figure 13 provides a copy of Figure 8 with both accuracy and NLL loss comparable for easy contrast. We note that high-performance clusters are distinguishable in both accuracy and loss, but the thresholded accuracy conceals low-performance clusters which are visible in the continuous distribution.
Figure 13: KDE approximations of performance variation on the MMLU test set. Provided for contrasting accuracy and continuous NLL loss. Top: Approximate distribution of accuracy metrics. Bottom: Approximate distribution of KDE metrics, copied from Figure 8.
@@
Figure 14: Histograms of the minimum probability of any token in each sequence, averaged across sequences (addition task). Metrics are calculated by Equation 2. The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Figure 4.
@@
Figure 16: Random variation in continuous length generalization error (count task). Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs. At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
Appendix E Additional Language Model Results
E.1 MMLU Results on Qwen2.5-1.5B
We repeat the same MMLU experiment (Section 3.1) on Qwen2.5-1.5B and we include results in Figure 17. Due to computation limit, for each data mixes, we train on 20 random seeds. Similar to Figure 13. models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
@@ Appendix E Additional Language Model Results
E.2 CSQA LM Results
In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA, Talmor et al. [2019]) task and we include the result in Figure 18. Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
Figure 18: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.

### v5

@@ Appendix B Breakthroughness and Linearity
Srivastava et al. (2023) introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})}$ (2)
$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$ (3)
@@
In Fig. 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for (a) count and (b) addition.
@@
Figure 10: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appx. B, with resulting seeds plotted for 10(a) count and 10(b) addition.
Appendix C Count task
Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training.
Zhou et al. (2024a) showed that models trained to count can generalize to more than twice this training length; however, Appx. C.2 reveals a more nuanced view of length generalization based on its distribution across independent model runs.
C.1 Count performance distributions
@@
Figure 11: Random variation in length generalization (count task). Histograms of exact match accuracy on length 6060 sequences when independently scaling (11(a)). width and (11(b)). depth.
@@
Our plots in the main paper show distributions of addition task accuracy with a fixed test length 40 (Fig. 2). In Fig. 11 we show analogous histograms for the count task with a fixed test length of 60. We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
In Sec. 2.5, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task. For example, Fig. 12 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
C.2 Count task as a case of U-shaped random scaling
We next consider the counting task (Sec. 2.1), which also exhibits bimodally-distributed performance (see Appx. Fig. 11) but yields a very different scaling effect: a U-shaped curve. Fig. 14 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Fig. 13.
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al. (2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Fig. 14 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 13: Changes in random variation (count task). Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently. We mark the emergence of bimodality at the last scale before multiple peaks appear. We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Fig. 14.
Figure 14: As in Fig. 3, we capture summary statistics for test length 60 on the count task. For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case. The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
@@ Appendix D Additional Continuous Metric Plots
Addition Task.
In Sec. 2.4 we have seen seen that the addition histograms are bimodal in exact match accuracy (Fig. 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Fig. 4 a, b ). We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
@@
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Fig.s 15 and 16 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
Count Task.
In Fig. 17, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Fig.s 11 and 16 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
Large LM Experiments.
Finally, Fig. 19 provides a copy of Fig. 9 with both accuracy and NLL loss comparable for easy contrast. We note that high-performance clusters are distinguishable in both accuracy and loss, but the thresholded accuracy conceals low-performance clusters which are visible in the continuous distribution. We also include histogram version of results in Fig. 19. Due to computation limit, for each data mixes, we train on 20 random seeds. Similar to Fig. 19. models yields the same bimodal distribution in accuracy metrics as well as multi-modal distribution in continuous metric (NLL).
@@
Figure 15: Histograms of the minimum probability of any token in each sequence, averaged across sequences (addition task). Metrics are calculated by Equation 4. The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Fig. 4 (a, b).
@@
Figure 17: Random variation in continuous length generalization error (count task). Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs. At scales where the EM accuracy distribution is most strongly bimodal (width=4, width=8, and depth=2), KDE places the most density at areas of high loss.
Figure 18: KDE approximations of performance variation on the MMLU test set. Provided for contrasting accuracy and continuous NLL loss. Top: Approximate distribution of accuracy metrics. Bottom: Approximate distribution of KDE metrics, copied from Fig. 9.
@@
Appendix E Grammatical Generalization Details
E.1 Task Description
In the question formation (QF) task, models transform declarative English sentences into questions by moving an auxiliary verb to the front of the sentence. The task tests whether models learn the correct hierarchical syntactic rule based on the sentence’s syntax tree structure, or instead rely on a superficial linear heuristic.
Our training data (based on McCoy et al. (2018)) permits two strategies for choosing which verb to move:
  1. 1.
    Linear rule: Move the first auxiliary verb in the sentence (incorrect for English)
  2. 2.
    Hierarchical rule: Move the main auxiliary verb, determined by the sentence’s syntactic structure (correct for English)
Examples of each rule are provided in Table 1. The first example is ambiguous because both the hierarchical and linear rules produce the same correct output. In contrast, the second example is unambiguous because only the hierarchical rule produces the grammatically correct question.
E.2 Data Composition
The training data contains two components:
  1. 1.
    Question formation examples: Transform declarative sentences to questions. These examples are always ambiguous between linear and hierarchical rules.
  2. 2.
    Declaration copying examples: Simply repeat declarative sentences without transformation. Unlike QF examples, these can include unambiguous sentence structures.
Following Qin et al. (2024), we manipulate the declaration copying subset to control exposure to different syntactic structures. Specifically, we vary the ratio of center-embedded sentences versus right-branching sentences:
  • Center-embedded sentences contain recursive clause embedding (e.g., “My unicorn who doesn’t sing does move”), which requires hierarchical syntactic processing and encourages learning the hierarchical rule.
  • Right-branching sentences have simpler linear structure (e.g., “My unicorn does move the dogs that do wait”), which can be solved with linear n-gram patterns.
We create training sets with different mixtures of these sentence types in the declaration copying subset: Quest Only (no declaration copying), Center Embed (only center-embedded declarations), Right Branch (only right-branching declarations), and various intermediate mixtures.
E.3 Evaluation
The in-distribution test set contains only ambiguous examples (solvable by both rules), while the OOD test set contains only unambiguous examples (requiring the hierarchical rule). Models using the hierarchical rule achieve $\sim$100% accuracy on both test sets. Models using the linear rule achieve $\sim$100% on in-distribution data but $\sim$0% on OOD data. We therefore use OOD accuracy as our measure of whether models have acquired the hierarchical syntactic rule.
Table 1: Question Formation task examples. Models must move the main auxiliary verb to the front to form a question. Ambiguous examples can be solved by either linear or hierarchical rules; unambiguous examples require the hierarchical rule.
Task Example Type Examples
Question Formation Ambiguous Input: My unicorn does move the dogs that do wait.
Output: Does my unicorn move the dogs that do wait?
Unambiguous Input: My unicorn who doesn’t sing does move.
Linear Output: Doesn’t my unicorn who sing does move?
Hierarchical Output: Does my unicorn who doesn’t sing move?